# Dynamic Injection of Entity Knowledge into Dense Retrievers

### v1

@@ Abstract
Dense retrievers often struggle with queries involving less-frequent entities due to their limited entity knowledge. We propose the Knowledgeable Passage Retriever (KPR), a BERT-based retriever enhanced with a context-entity attention layer and dynamically updatable entity embeddings. This design enables KPR to incorporate external entity knowledge without retraining. Experiments on three datasets show that KPR consistently improves retrieval accuracy, achieving a substantial 12.6% gain on the EntityQuestions dataset over the model without KPR extensions. When built on the off-the-shelf bge-base retriever, KPR achieves state-of-the-art performance among similarly sized models on two datasets. Code and models will be released soon.
@@ 1 Introduction
In this paper, we address this problem by proposing a simple extension to dense retrievers that dynamically injects entity knowledge into their embeddings. Specifically, we introduce the Knowledgeable Passage Retriever (KPR), a BERT-based dense retriever Karpukhin et al. (2020) that integrates entity embeddings adaptable at inference time (see Figure 1). KPR is trained to attend to entity knowledge based on the context of the input text, through a context-entity attention layer placed on top of BERT. Entity embeddings are obtained via single-pass BERT inference on texts referring to the entity and are kept frozen during training. KPR detects entities in the input text using a simple dictionary-based entity linker. Since both the entity embeddings and the linker can be modified after training, new entity knowledge can be easily added or updated without retraining.
@@ 2 Method
DPR. DPR encodes a query or passage into a -dimensional embedding, obtained from the output embedding of BERT’s [CLS] token. Relevance between a query and a passage is scored as the dot product of their embeddings. The model is fine-tuned on datasets containing queries paired with positive and negative passages, using a cross-entropy loss with the dot product scores as logits. See Appendix A for further details.
Input entity representation. KPR uses the following two -dimensional input embeddings:
@@
Context-entity attention layer. KPR adopts a single-head key-query-value attention mechanism Vaswani et al. (2017). The query matrix is computed based on the output embedding of BERT’s [CLS] token (denoted by ). The key and value matrices, and , are computed based on the input entity representations denoted by :
@@
Entity embedder. We compute entity embeddings based on BERT inference over Wikipedia passages that refer to the corresponding entity Ye et al. (2022). Specifically, for each passage containing an anchor link to the entity, we replace the entity name with BERT’s [MASK] token, extract the output embedding of the [MASK] token, and average the resulting embeddings across passages. The final embeddings are then normalized to match the average norm of BERT’s input token embeddings.
Entity linker. We use dictionary-based string matching to detect entities mentioned in the text. The dictionary comprises anchor names and their possible referent entities, obtained from Wikipedia hyperlinks. For ambiguous names (e.g., “Apple,” which can refer to Apple Inc. and Apple (food)), we do not disambiguate to a single entity, but instead input all possible candidates to maintain recall.
Training. The parameters in our attention layer, entity position embeddings, and no-op embedding are initialized randomly. Entity embeddings are kept frozen during training, enabling them to be added or updated dynamically after training. The dropout probability in the attention layer is set to match that of the other layers. We follow DPR Karpukhin et al. (2020) for the remaining training settings including the loss function. Note that entities can be dynamically updated in KPR without retraining, since KPR relies on entity embeddings computed via BERT inference, which are kept frozen during training, and on a dictionary-based entity linker, whose entries can be modified.
Computational overhead. KPR adds entity embeddings to BERT, requiring additional storage. However, since the embeddings are used only as input features, they can be implemented as a sparse lookup table and stored not only in GPU memory but also in CPU memory or even on disk, with minimal impact on speed Yu et al. (2025). Furthermore, KPR introduces only a small overhead in total FLOPs, as discussed in Appendix B.
@@ 3 Experiments
Entity set. We use 7.2M English Wikipedia entities to construct our entity embeddings and linker.
Datasets and metrics. We train our model on the dataset proposed by Karpukhin et al. (2020), which is based on NQ, TQA, WebQuestions Berant et al. (2013), and CuratedTREC Baudiš and Šedivý (2015). We evaluate the models on the EQ, NQ, and TQA datasets. Following Sciavolino et al. (2021), we report top 20 retrieval accuracy. Top 100 accuracies are also provided in Appendix D. We use the 21M Wikipedia passages released by Karpukhin et al. (2020) as the target passages.
Baselines. We use vanilla DPR and conventional BM25, which performs robustly on queries with less frequent entities Sciavolino et al. (2021).
Base models. We use base-sized BERT and RetroMAE Xiao et al. (2022), a state-of-the-art BERT-based model pretrained for retrieval. For DPR, we also adopt PELT Ye et al. (2022), a BERT-based model that injects entity knowledge by inserting each entity embedding, enclosed between the token embeddings of parentheses, immediately after the entity’s token embeddings. PELT introduces entity knowledge via input augmentation, whereas KPR incorporates it through a dedicated attention layer.
Entity embeddings. We use the same entity embeddings, derived from base-sized BERT inference (§2), for both PELT and all KPR models.
Entity linker. By default, we use our dictionary-based linker (§2) and additionally employ the state-of-the-art ReFinED linker Ayoola et al. (2022).
@@ 3.1 Results and Analysis
Model | Base Model | Entity Linker | EQ | NQ | TQA | Avg |
BM25 | – | – | 71.2 | 59.1 | 66.9 | 66.0 |
DPR | BERT | – | 56.7 | 79.4 | 78.8 | 71.6 |
KPR | BERT | Dictionary | 69.3 | 80.6 | 79.7 | 76.5 |
DPR | BERT | ReFinED | 64.0 | 79.9 | 78.9 | 74.3 |
KPR | BERT | ReFinED | 68.8 | 80.3 | 79.8 | 76.3 |
DPR | RetroMAE | – | 71.0 | 81.1 | 81.1 | 77.7 |
KPR | RetroMAE | Dictionary | 74.6 | 82.9 | 82.1 | 79.9 |
@@
Sigmoid function with length bias (KPR) | 69.3 | 80.6 | 79.7 | 76.5 |
@@
Main results. Table 1 shows that KPR consistently outperforms all baselines across datasets and base models. KPR surpasses DPR by 12.6% on EQ and 4.9% on average. Using the same ReFinED linker, KPR consistently outperforms DPR, with a 4.8% gain on EQ, suggesting that the entity knowledge injected by PELT is partially lost within BERT and highlighting the benefit of KPR’s attention layer placed on top. KPR achieves the best performance across all datasets, demonstrating its effectiveness even with a strong pretrained model. The gain is notably larger on EQ in all settings, as it includes many queries involving less-frequent entities. Further analysis of the EQ results is provided below.
@@
Effects of entity linker. Table 1 also shows that KPR with the dictionary-based linker outperforms its ReFinED-based variant on average. This is somewhat surprising, as the dictionary linker simply applies string matching without disambiguation (§2) and may detect incorrect or noisy entities. We attribute this to ReFinED’s slightly lower recall, as it extracts entities only when confident. For example, ReFinED detects 0.93 entities per query on average in EQ, compared to 0.97 with the dictionary linker. These results also suggest that KPR is robust to noise, likely due to its attention mechanism’s ability to focus on contextually relevant entities. See Appendix F for further analysis.
How do entity embeddings affect performance? We evaluate two baselines for computing entity embeddings: random initialization and Wikipedia2Vec embeddings Yamada et al. (2020a), both based on the same Wikipedia dump as our entity embeddings. We also test embeddings extracted from an intermediate BERT layer, motivated by prior work suggesting that entity knowledge is well captured there Meng et al. (2022).
@@
Sigmoid vs. softmax. We evaluate the effect of replacing the sigmoid activation in our attention layer with the softmax function, which is commonly used in attention mechanisms. Table 2 shows that sigmoid consistently outperforms softmax across all datasets. We attribute this to the fact that sigmoid allows the model to assess each entity’s relevance independently, without being influenced by the presence of other relevant entities.
Analysis of EQ results. To examine whether KPR’s improvement stems from incorporating knowledge of less-frequent entities, we divide the EQ examples into 10 bins based on the frequency of the entity in each query and measure performance within each bin. This is feasible because every EQ query contains a single entity. We obtain entity frequencies from Wikipedia hyperlinks and create 10 log-spaced bins ranging from 1 to 10,000.
@@
A detailed analysis of the behavior of our attention mechanism is presented in Appendix F.
@@ Appendix A Overview of DPR
Training. Let be a set of training instances, where each instance consists of a query , a positive passage relevant to the query, and negative passages irrelevant to the query. The model is trained by minimizing the negative log-likelihood of the positive passage:
@@
Inference. DPR constructs a passage index by encoding all target passages. At runtime, it retrieves the top-ranked passages using maximum inner product search, with the query embedding as input.
@@ Appendix B Notes on Computational Overhead
Top 20 | Top 100 | |||||||||
Model | Base Model | Entity Linker | EQ | NQ | TQA | Avg | EQ | NQ | TQA | Avg |
BM25 | – | – | 71.2 | 59.1 | 66.9 | 66.0 | 79.8 | 73.7 | 76.7 | 76.7 |
DPR | BERT | – | 56.7 | 79.4 | 78.8 | 71.6 | 70.0 | 86.0 | 84.7 | 80.2 |
KPR | BERT | Dictionary | 69.3 | 80.6 | 79.7 | 76.5 | 78.9 | 87.0 | 85.0 | 83.6 |
DPR | BERT | ReFinED | 64.0 | 79.9 | 78.9 | 74.3 | 75.3 | 86.3 | 84.7 | 82.1 |
KPR | BERT | ReFinED | 68.8 | 80.3 | 79.8 | 76.3 | 78.7 | 86.7 | 85.0 | 83.5 |
DPR | RetroMAE | – | 71.0 | 81.1 | 81.1 | 77.7 | 79.8 | 87.7 | 86.0 | 84.5 |
KPR | RetroMAE | Dictionary | 74.6 | 82.9 | 82.1 | 79.9 | 82.3 | 88.2 | 86.7 | 85.7 |
Top 20 | Top 100 | |||||||
Model | EQ | NQ | TQA | Avg | EQ | NQ | TQA | Avg |
contriever Izacard et al. (2022) | 63.0 | 67.9 | 73.9 | 68.3 | 75.2 | 80.6 | 82.9 | 80.0 |
gte-base Li et al. (2023) | 70.8 | 77.0 | 76.6 | 74.7 | 79.8 | 86.4 | 83.9 | 83.4 |
e5-base Wang et al. (2022) | 72.4 | 86.2 | 81.4 | 80.0 | 82.5 | 90.8 | 86.2 | 86.5 |
bge-base Xiao et al. (2024) | 71.0 | 82.3 | 80.0 | 77.8 | 80.8 | 88.5 | 85.9 | 85.1 |
KPR | 76.8 | 82.4 | 81.5 | 80.2 | 84.2 | 89.0 | 86.6 | 86.6 |
@@ Appendix C Detailed Experimental Setup
Entity vocabulary. The entity vocabulary of KPR consists of 7.2M English Wikipedia entities that appear as hyperlinks at least once in the April 2024 dump. These hyperlinks are extracted using the mwparserfromhell library.111https://github.com/earwig/mwparserfromhell
Entity embeddings. As described in Section 2, entity embeddings are obtained by running BERT inference on Wikipedia passages that refer to the corresponding entities. For each entity, we randomly select up to 128 such passages and compute the embedding using the method outlined in Section 2.
@@
Entity linker. The dictionary used in our entity linker is constructed directly from entity hyperlinks in Wikipedia. For example, if a hyperlink with the anchor text “New York” refers to the entity New York City, we register “New York” as an entity name and New York City as its possible referent.
@@
Model. We use the base-sized BERT333https://huggingface.co/google-bert/bert-base-uncased with 110M parameters. RetroMAE444https://huggingface.co/Shitao/RetroMAE and bge-base555https://huggingface.co/BAAI/bge-base-en are also based on this BERT model. KPR with the entity embeddings contains 5.6B parameters.
Datasets. The licenses and the number of training, development, and test examples for the datasets used in this paper are provided in Table 4. All datasets are publicly and freely available for research purposes. Access to CuratedTREC is governed by the TREC organizers. MS MARCO is freely available for non-commercial research purposes.666https://microsoft.github.io/msmarco/ We also use Wikipedia, which is licensed under CC BY-SA 4.0.
@@
Training. Our training settings follow Karpukhin et al. (2020). For each query, we use one positive and one hard negative passage, and construct a mini-batch of 128 queries. We share parameters between the query and passage encoders to reduce computational cost, as using separate encoders yields nearly identical results (within 0.2% difference in top-20 accuracy across datasets) in our experiments, consistent with prior work Lee et al. (2023); Zhang et al. (2023a). We optimize the model using Adam Kingma and Ba (2014) with a learning rate of 2e-5 for 40 epochs. We apply in-batch negatives (see Appendix A) during training. We do not perform hyperparameter tuning and instead use the same settings provided in the GitHub repository of Karpukhin et al. (2020).777https://github.com/facebookresearch/DPR/blob/main/conf/train/biencoder_nq.yaml
@@
Appendix D Full Experimental Results

### v2

@@ Abstract
Dense retrievers often struggle with queries involving less-frequent entities due to their limited entity knowledge. We propose the Knowledgeable Passage Retriever (KPR), a BERT-based retriever enhanced with a context-entity attention layer and dynamically updatable entity embeddings. This design enables KPR to incorporate external entity knowledge without retraining. Experiments on three datasets demonstrate that KPR consistently improves retrieval accuracy, with particularly large gains on the EntityQuestions dataset. When built on the off-the-shelf bge-base retriever, KPR achieves state-of-the-art performance among similarly sized models on two datasets. Models and code are released at github.com/knowledgeable-embedding/knowledgeable-embedding.
@@ 1 Introduction
In this paper, we address this problem by proposing a simple extension to dense retrievers that dynamically injects entity knowledge into their embeddings. Specifically, we introduce the Knowledgeable Passage Retriever (KPR), a BERT-based dense retriever Karpukhin et al. (2020) that integrates entity embeddings adaptable at inference time (see Figure 1). KPR is intentionally designed with a simple architecture and is trained to attend to entity knowledge based on the context of the input text, through a context-entity attention layer placed on top of BERT. Entity embeddings are obtained via single-pass BERT inference on texts referring to the entity and are kept frozen during training. KPR detects entities in the input text using a simple dictionary-based entity linker. Since both the entity embeddings and the linker can be modified after training, new entity knowledge can be easily added or updated without retraining.
@@ 2 Method
DPR.
DPR encodes a query or passage into a -dimensional embedding, obtained from the output embedding of BERT’s [CLS] token. Relevance between a query and a passage is scored as the dot product of their embeddings. The model is fine-tuned on datasets containing queries paired with positive and negative passages, using a cross-entropy loss with the dot product scores as logits. See Appendix A for further details.
Input entity representation.
KPR uses the following two -dimensional input embeddings:
@@
Context-entity attention layer.
KPR adopts a single-head key-query-value attention mechanism Vaswani et al. (2017). The query matrix is computed based on the output embedding of BERT’s [CLS] token (denoted by ). The key and value matrices, and , are computed based on the input entity representations denoted by :
@@
Entity embedder.
We compute entity embeddings based on BERT inference over Wikipedia passages that refer to the corresponding entity Ye et al. (2022). Specifically, for each passage containing an anchor link to the entity, we replace the entity name with BERT’s [MASK] token, extract the output embedding of the [MASK] token, and average the resulting embeddings across passages. The final embeddings are then normalized to match the average norm of BERT’s input token embeddings.
Entity linker.
We use dictionary-based string matching to detect entities mentioned in the text. The dictionary comprises anchor names and their possible referent entities, obtained from Wikipedia hyperlinks. For ambiguous names (e.g., “Apple,” which can refer to Apple Inc. and Apple (food)), we do not disambiguate to a single entity, but instead input all possible candidates to maintain recall.
Training.
The parameters in our attention layer, entity position embeddings, and no-op embedding are initialized randomly. Entity embeddings are kept frozen during training, enabling them to be added or updated dynamically after training. The dropout probability in the attention layer is set to match that of the other layers. We follow DPR Karpukhin et al. (2020) for the remaining training settings including the loss function. Note that entities can be dynamically updated in KPR without retraining, since KPR relies on entity embeddings computed via BERT inference, which are kept frozen during training, and on a dictionary-based entity linker, whose entries can be modified.
Computational overhead.
KPR adds entity embeddings to BERT, requiring additional storage. However, since the embeddings are used only as input features, they can be implemented as a sparse lookup table and stored not only in GPU memory but also in CPU memory or even on disk, with minimal impact on speed Yu et al. (2025). Furthermore, KPR introduces only a small overhead in total FLOPs, as discussed in Appendix B.
@@ 3 Experiments
Entity set.
We use 7.2M English Wikipedia entities to construct our entity embeddings and linker.
Datasets and metrics.
We train our model on the dataset proposed by Karpukhin et al. (2020), which is based on NQ, TQA, WebQuestions Berant et al. (2013), and CuratedTREC Baudiš and Šedivý (2015). We evaluate the models on the EQ, NQ, and TQA datasets. Following Sciavolino et al. (2021), we report top 20 retrieval accuracy. Top 100 accuracies are also provided in Appendix D. We use the 21M Wikipedia passages released by Karpukhin et al. (2020) as the target passages.
Baselines.
We use vanilla DPR and conventional BM25, which performs robustly on queries with less frequent entities Sciavolino et al. (2021).
Base models.
We use base-sized BERT and RetroMAE Xiao et al. (2022), a state-of-the-art BERT-based model pretrained for retrieval. For DPR, we also adopt PELT Ye et al. (2022), a BERT-based model that injects entity knowledge by inserting each entity embedding, placed between the token embeddings of parentheses, immediately after the corresponding entity token. While PELT introduces entity knowledge through input augmentation, KPR incorporates it via a dedicated attention layer. Unlike KPR, PELT requires each input entity name to be disambiguated to a single entity, which incurs additional computational complexity.
Entity embeddings.
We use the same entity embeddings, derived from base-sized BERT inference (§2), for both PELT and all KPR models.
Entity linker.
By default, we use our dictionary-based linker (§2) and additionally employ the state-of-the-art ReFinED linker Ayoola et al. (2022). We select ReFinED over other off-the-shelf systems such as BLINK Wu et al. (2020), GENRE Cao et al. (2021), and LUKE Yamada et al. (2022) due to its superior performance.
@@ 3.1 Results and Analysis
| Model | Base Model | Entity Linker | EQ | NQ | TQA | Avg | ||||||
| BM25 | – | – | 71.2 | 59.1 | 66.9 | 66.0 | ||||||
| DPR | BERT | – |
71.7 | ||||||
| KPR | BERT | Dictionary |
76.6 | ||||||
| DPR | BERT | ReFinED |
74.2 | ||||||
| KPR | BERT | ReFinED |
76.4 | ||||||
| DPR | RetroMAE | – |
77.7 | ||||||
| KPR | RetroMAE | Dictionary |
79.9 |
@@
| Sigmoid function with length bias | 69.3 | 80.6 | 79.7 | 76.5 |
@@
Main results.
Table 1 shows that KPR significantly outperforms all baselines across datasets and base models. KPR surpasses DPR by 12.6% on EQ and 4.9% on average. Using the same ReFinED linker, KPR consistently outperforms DPR, with a 5.2% gain on EQ, suggesting that the entity knowledge injected by PELT is partially lost within BERT and highlighting the benefit of KPR’s attention layer placed on top. KPR achieves the best performance across all datasets, demonstrating its effectiveness even with a strong pretrained model. The gain is notably larger on EQ in all settings, as it includes many queries involving less-frequent entities. Further analysis of the EQ results is provided below.
@@
Effects of entity linker.
Table 1 also shows that KPR with the dictionary-based linker outperforms its ReFinED-based variant on average. This is somewhat surprising, as the dictionary linker simply applies string matching without disambiguation (§2) and may detect incorrect or noisy entities. We attribute this to ReFinED’s slightly lower recall, as it extracts entities only when confident. For example, ReFinED detects 0.93 entities per query on average in EQ, compared to 0.97 with the dictionary linker. These results also suggest that KPR is robust to noise, likely due to its attention mechanism’s ability to focus on contextually relevant entities. See Appendix F for further analysis.
How do entity embeddings affect performance?
We evaluate two baselines for computing entity embeddings: random initialization and Wikipedia2Vec embeddings Yamada et al. (2020a), both based on the same Wikipedia dump as our entity embeddings. We also test embeddings extracted from an intermediate BERT layer, motivated by prior work suggesting that entity knowledge is well captured there Meng et al. (2022).
@@
Sigmoid vs. softmax.
We evaluate the effect of replacing the sigmoid activation in our attention layer with the softmax function, which is commonly used in attention mechanisms. Table 2 shows that sigmoid consistently outperforms softmax across all datasets. We attribute this to the fact that sigmoid allows the model to assess each entity’s relevance independently, without being influenced by the presence of other relevant entities.
Analysis of EQ results.
To examine whether KPR’s improvement stems from incorporating knowledge of less-frequent entities, we divide the EQ examples into 10 bins based on the frequency of the entity in each query and measure performance within each bin. This is feasible because every EQ query contains a single entity. We obtain entity frequencies from Wikipedia hyperlinks and create 10 log-spaced bins ranging from 1 to 10,000.
@@
Analysis of KPR’s Attention Mechanism.
@@ Appendix A Overview of DPR
Training.
Let be a set of training instances, where each instance consists of a query , a positive passage relevant to the query, and negative passages irrelevant to the query. The model is trained by minimizing the negative log-likelihood of the positive passage:
@@
Inference.
DPR constructs a passage index by encoding all target passages. At runtime, it retrieves the top-ranked passages using maximum inner product search, with the query embedding as input.
@@ Appendix B Notes on Computational Overhead
| Model | Base Model | Entity Linker | EQ | NQ | TQA | Avg |
| BM25 | – | – | 79.8 | 73.7 | 76.7 | 76.7 |
| DPR | BERT | – | 70.0 | 86.0 | 84.7 | 80.2 |
| KPR | BERT | Dictionary | 78.9 | 87.0 | 85.0 | 83.6 |
| DPR | BERT | ReFinED | 75.3 | 86.3 | 84.7 | 82.1 |
| KPR | BERT | ReFinED | 78.7 | 86.7 | 85.0 | 83.5 |
| DPR | RetroMAE | – | 79.8 | 87.7 | 86.0 | 84.5 |
| KPR | RetroMAE | Dictionary | 82.3 | 88.2 | 86.7 | 85.7 |
| Model | EQ | NQ | TQA | Avg |
| contriever Izacard et al. (2022) | 75.2 | 80.6 | 82.9 | 80.0 |
| gte-base Li et al. (2023) | 79.8 | 86.4 | 83.9 | 83.4 |
| e5-base Wang et al. (2022) | 82.5 | 90.8 | 86.2 | 86.5 |
| bge-base Xiao et al. (2024) | 80.8 | 88.5 | 85.9 | 85.1 |
| KPR | 84.2 | 89.0 | 86.6 | 86.6 |
@@ Appendix C Detailed Experimental Setup
Entity vocabulary.
The entity vocabulary of KPR consists of 7.2M English Wikipedia entities that appear as hyperlinks at least once in the April 2024 dump. These hyperlinks are extracted using the mwparserfromhell library.111https://github.com/earwig/mwparserfromhell
Entity embeddings.
As described in Section 2, entity embeddings are obtained by running BERT inference on Wikipedia passages that refer to the corresponding entities. For each entity, we randomly select up to 128 such passages and compute the embedding using the method outlined in Section 2.
@@
Entity linker.
The dictionary used in our entity linker is constructed directly from entity hyperlinks in Wikipedia. For example, if a hyperlink with the anchor text “New York” refers to the entity New York City, we register “New York” as an entity name and New York City as its possible referent.
@@
Model.
We use the base-sized BERT333https://huggingface.co/google-bert/bert-base-uncased with 110M parameters. RetroMAE444https://huggingface.co/Shitao/RetroMAE and bge-base555https://huggingface.co/BAAI/bge-base-en are also based on this BERT model. KPR with the entity embeddings contains 5.6B parameters.
Datasets.
The licenses and the number of training, development, and test examples for the datasets used in this paper are provided in Table 4. All datasets are publicly and freely available for research purposes. Access to CuratedTREC is governed by the TREC organizers. MS MARCO is freely available for non-commercial research purposes.666https://microsoft.github.io/msmarco/ We also use Wikipedia, which is licensed under CC BY-SA 4.0.
@@
Training.
Our training settings follow Karpukhin et al. (2020). For each query, we use one positive and one hard negative passage, and construct a mini-batch of 128 queries. We share parameters between the query and passage encoders to reduce computational cost, as using separate encoders yields nearly identical results (within 0.2% difference in top 20 accuracy across datasets) in our experiments, consistent with prior work Lee et al. (2023); Zhang et al. (2023a). We optimize the model using Adam Kingma and Ba (2014) with a learning rate of 2e-5 for 40 epochs. We apply in-batch negatives (see Appendix A) during training. We do not perform hyperparameter tuning and instead use the same settings provided in the GitHub repository of Karpukhin et al. (2020).777https://github.com/facebookresearch/DPR/blob/main/conf/train/biencoder_nq.yaml
@@
Appendix D Results Based on Top 100 Accuracy